OpenStudy (anonymous):

The label on a bottle reads "MOH" and you set out to find the identity of the metal (M). Assuming that this is an alkali metal hydroxide, you weigh out 1.67 grams of the solid. It takes 0.0418L of 1.10M HCl to titrate the metal hydroxide. What is the identity of the solid?

OpenStudy (anonymous):

Please show all calc involved

OpenStudy (anonymous):

Using C = n/v You can use that to find the number of mols of HCl required to neutralize the hydroxide compound. Once you've found the number of mols of HCl, there is a 1:1 ratio between H+ and OH-, knowing the fact that "M" is an alkali metal, ( 1 M+ for each OH-) So therefore the number of mols of HCl used is equal to the number of mols of MOH used, so knowing the mols, you can find its total molar mass n = m / MM, where n is the number of mols, m is the mass, and MM is the molar mass After finding the total molar mass you can subtract off the "OH" molar masses, to find the identity of M

OpenStudy (anonymous):

Equation: MOH + HCl → MCl + H2O 1mol MOH reacts with 1mol HCl Mol HCl in 0.0418L of 1.10M HCl = 0.0418*1.10 = 0.04598 mol HCl Therefore you have 0.04598 mol MOH 0.04598 mol MOH has mass = 1.67g 1mol MOH has mass 1.67/0.04598 = 36.32g/mol when i subtract 17.00(molar mass of OH) from 36.32g/mol. I get 19.36. It doesnt equal to any of the amu of the alkali metal. where did I go wrong?

OpenStudy (anonymous):

I'm pretty sure the process is correct, and I also did get the same answer. If the molar concentration was instead 1.01M HCl, we'd get a closer answer of around ~~ 22.5g/mol, that of sodium.